# Finetuning Large Language Model as an Effective Symbolic Regressor

### v1

@@ 3. Final Score Aggregation
$S_{\text{struct}}=\frac{1}{6}\sum_{k=1}^{6}\text{Sim}_{k},$ (11)
@@
More Experiments and Visualizations
GRPO-hyperprameter Selection
In this section, we conduct an ablation study on the key hyperparameters of our proposed GRPO method. The results are summarized in Table 8.
Based on the ablation study in Table 5, we selected configuration (f) for our SymbolicChat. Although other settings like (c) yield higher structural similarity, configuration (f) achieves the highest numerical accuracy representing the best compromise between finding a structurally sound equation and achieving a precise data fit.
Results visualization
@@
To provide a qualitative assessment of our method’s performance, we present a case study in Table 6. The objective is to recover the ground-truth equation $C*x_{1}+C*arctan(C*x_{0}+C)+C$. As shown, our proposed method, SymbolicChat, is the only one capable of precisely identifying the exact symbolic form of the ground-truth equation.
@@
SymbolicChat(Ours) $C*x_{1}+C*arctan(C*x_{0}+C)+C$
@@
Table 6: Comparison with Baseline Methods.
@@
Table 7: Example of results of SymbolicChat
(a) 1.0 1.0 1.0 1.0 0.416 0.595 0.516 0.260
(b) 1.0 1.0 2.0 1.0 0.403 0.576 0.539 0.260
(c) 1.0 2.0 1.0 1.0 0.431 0.604 0.490 0.256
(d) 1.0 2.0 2.0 1.0 0.418 0.597 0.521 0.256
(e) 1.0 2.0 2.0 2.0 0.418 0.594 0.527 0.244
(f) 1.0 2.0 2.0 4.0 0.403 0.576 0.540 0.252
Table 8: Impact of reward hyperparameter settings on model performance. Each row represents a different configuration of reward weights.

### v2

@@ 3. Final Score Aggregation
$Sim=\frac{1}{6}\sum_{k=1}^{6}\text{Sim}_{k},$ (11)
@@
Appendix C More Experiments and Visualizations
C.1 Ablation Study on Form-GRPO Hyperparameter
To determine the optimal composition of our reward function, we conducted a series of experiments to analyze the impact of different reward hyperparameter weights on model performance. As detailed in Tab. 6, we evaluated six distinct configurations by varying the weights for form similarity reward(R_similarity), numerical reward (R_numerical), and equivalent reward (R_equiv). The form-correct reward weight (R_format) was held constant at 1.0 across all experiments, as we consider format correctness a foundational requirement.
Table 6: Ablation study on Form-GRPO hyperparameter settings on model performance. Each row represents a different configuration of reward weights.
(a) 1.0 1.0 1.0 1.0 0.431 0.598 0.751 0.379
(b) 1.0 1.0 2.0 1.0 0.405 0.580 0.814 0.412
(c) 1.0 2.0 1.0 1.0 0.435 0.606 0.738 0.375
(d) 1.0 2.0 2.0 1.0 0.435 0.605 0.794 0.391
(e) 1.0 2.0 2.0 2.0 0.433 0.607 0.797 0.392
(f) 1.0 2.0 2.0 4.0 0.436 0.607 0.808 0.404
Our analysis began with a baseline configuration (a), where all reward components were equally weighted, yielding an $R^{2}$ score of 0.751. While increasing only the numerical reward weight (configuration b) resulted in the highest $R^{2}$ (0.814) and $\text{Acc}_{r}$ (0.412), we observed that this came at the cost of structural quality, as indicated by lower $\text{S}_{\text{struct}}$ scores.
Our goal was to find a configuration that achieves a robust balance across all performance dimensions. Through systematic adjustments, we identified configuration (f), with weights of 1.0, 2.0, 2.0, and 4.0 for R_format, R_similarity, R_numerical, and R_equiv respectively. This configuration achieved the highest scores for structural similarity ($\text{S}_{\text{struct}}\text{(gpt-4o)}=0.436$ and $\text{S}_{\text{struct}}\text{(rule)}=0.607$) among all tested setups.
Although configuration (f) exhibits a marginal decrease in $R^{2}$ (0.808) and $\text{Acc}_{r}$ (0.404) compared to the peak values in configuration (b), its superior performance in structural metrics demonstrates a more desirable trade-off. The strong emphasis on equivalent reward (R_equiv = 4.0) proved crucial for enhancing the model’s ability to generate outputs that are not only numerically accurate but also logically and structurally sound.
Therefore, we concluded that configuration (f) represents the optimal balance for our objectives. These hyperparameter weights were adopted for all main experiments.
C.2 Evaluating the Impact of Noise on Model Performance
To evaluate the model’s robustness against measurement uncertainties, we introduce additive zero-mean Gaussian noise to the output values ($y_{i}$) of the test set. The standard deviation of the noise is fixed at $\sigma=0.001$. This procedure is designed to simulate a low level of absolute measurement error and examine the model’s stability under this specific condition.
Table 7: Evaluation of model robustness to additive Gaussian noise ($\sigma=0.001$) on the SymbArena dataset. The best and second-best results are highlighted in bold and underlined, respectively.
Methods Type $S_{\text{struct}}$ (gpt-4o) $S_{\text{struct}}$ (rule) $R^{2}$ $\text{Acc}_{\tau}$
gplearn GP 0.261 0.301 0.209 0.074
AFP GP 0.293 0.396 0.276 0.076
AFP-FE GP 0.297 0.394 0.246 0.058
GP-GOMEA GP 0.337 0.362 0.334 0.214
PySR GP 0.371 0.377 0.625 0.341
LLM-SR(gpt-4o-mini) LLM-based 0.343 0.499 0.251 0.169
LLM-SR(Qwen2.5-7B) LLM-based 0.340 0.452 0.250 0.159
SGA(gpt-4o-mini) LLM-based 0.342 0.505 0.271 0.193
SGA(Qwen2.5-7B) LLM-based 0.352 0.488 0.253 0.175
Symbolic-R1 LLM-based 0.403 0.596 0.793 0.353
As presented in Tab. 7, the introduction of noise led to a predictable performance degradation across most baseline methods. Other competitive methods such as PySR experienced a more pronounced performance drop (e.g., its $R^{2}$ score fell from 0.663 to 0.625). In contrast, Symbolic-R1 demonstrated exceptional stability. Specifically, its performance exhibited only a marginal decline: the $R^{2}$ score decreased minimally from 0.808 to 0.793, and the rule-based structural similarity, $S_{\text{struct}}(\text{rule})$, dropped from 0.607 to just 0.596. This minimal decay underscores the model’s high resilience to low-level data perturbations.
C.3 Evaluation on More Datasets
To further assess the generalization capability of our proposed method, its performance was also evaluated on a set of five well-established benchmarks, including Nguyen (Uy et al., 2011), Constant (Tian et al., 2025), Keijzer (Keijzer, 2003), R rational (Krawiec & Pawlak, 2013), and SRBench (La Cava et al., 2021).
Table 8: Performance comparison on five well-established benchmarks. The best and second-best results are highlighted in bold and underlined, respectively.
Method Nguyen Constant R Keijzer SRBench Overall avg.
gplearn $0.767$ $0.606$ $0.354$ $0.638$ $0.551$ $0.621$
AFP $0.702$ $0.720$ $0.657$ $0.800$ $0.755$ $0.742$
AFP-FE $0.832$ $0.790$ $0.457$ $0.783$ $0.619$ $0.727$
GP-GOMEA $0.597$ $0.675$ $0.336$ $0.261$ $0.684$ $0.539$
PySR µ l 0.950 $0.998$ $0.997$ $0.999$ µ l 0.937 µ l 0.968
LLM-SR(gpt-4o-mini) $0.567$ $0.620$ $0.051$ $0.523$ $0.234$ $0.434$
LLM-SR(Qwen2.5-7B) $0.623$ $0.734$ $0.315$ $0.478$ $0.353$ $0.506$
SGA(gpt-4o-mini) $0.747$ $0.667$ $0.875$ $0.577$ $0.421$ $0.603$
SGA(Qwen2.5-7B) $0.535$ $0.686$ $0.611$ $0.465$ $0.286$ $0.472$
Symbolic-R1 (Ours) $0.973$ µ l 0.977 $0.997$ µ l 0.972 $0.955$ $0.969$
As presented in Tab. 8, our proposed method, Symbolic-R1, demonstrates state-of-the-art performance and robust generalization across five well-established benchmarks. It achieves the highest overall average score of 0.969, outperforming all competing methods. Specifically, Symbolic-R1 secures the top rank on three of the five benchmarks (Nguyen, R, and SRBench) and the second-best rank on the remaining two (Constant and Keijzer), consistently placing it at the forefront of performance. This strong result, which surpasses various baselines including recent LLM-based methods and is highly competitive with the powerful PySR, underscores the superior effectiveness and generalization capability of our approach.
C.4 Results visualization
@@
To provide a qualitative assessment of our method’s performance, we present a case study in Table 9. The objective is to recover the ground-truth equation $C*x_{1}+C*arctan(C*x_{0}+C)+C$. As shown, our proposed method, Symbolic-R1, is the only one capable of precisely identifying the exact symbolic form of the ground-truth equation.
@@
Table 9: A Visual Comparison with Baseline Methods.
@@
Symbolic-R1(Ours) $C*x_{1}+C*arctan(C*x_{0}+C)+C$
@@
Table 10: Example of results of Symbolic-R1